OpenStudy (anonymous):

Solve x3 = 64 over 27. ±8 over 3 8 over 3 ±4 over 3 4 over 3

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you could check your answer choices

OpenStudy (anonymous):

this isnt the one that needs checking

OpenStudy (phi):

x^3 means x*x*x so you could check if the answer times itself 3 times gives 64/27

OpenStudy (phi):

I am suggesting a way to decide what the answer is. one way is to check each of the 4 possible answers. only one of them should "work"

OpenStudy (phi):

the other way to answer the question is to "solve for x" but that involves finding cube roots.

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

let me look rlly fast

OpenStudy (anonymous):

im not sure

OpenStudy (phi):

let's test x= 8/3 what is \[ \frac{8}{3} \cdot \frac{8}{3} \cdot \frac{8}{3} \] multiply top times top and bottom times bottom

OpenStudy (anonymous):

512/27?

OpenStudy (anonymous):

@phi is that it

OpenStudy (phi):

yes and 512/27 is not 64/27 in other words, 8/3 (and also -8/3) is not the answer the other two choices are 4/3 and \(\pm\) 4/3 the only difference is the second choice includes -4/3 but -4/3 * -4/3 * -4/3= -64/27 it is negative and we want +64/27 that means the only answer that works is 4/3